OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (ybarrap):

Is that $$ 7\sin^2x-14\sin x+ 2=-5 $$ ?

OpenStudy (anonymous):

yes

OpenStudy (ybarrap):

set $$ y=\sin x $$ Then solve $$ 7y^2-14y+ 7=0 $$ Can you do that?

OpenStudy (ybarrap):

You can use the quadratic formula for that Once you get \(y\) Then solve for \(x\) $$ y=\sin x\\ x=\sin^{-1}y $$

OpenStudy (anonymous):

whats the answer?

OpenStudy (ybarrap):

It's your turn. Do you know the quadratic formula?